OpenStudy (anonymous):

find the absolute maximum and absolute minimum values of f on the given interval f(x)=ln(x^2+x+1) [-1,1]

OpenStudy (anonymous):

calculus or graph?

OpenStudy (anonymous):

Uhm, calculus.

OpenStudy (anonymous):

ok i guess we can take the derivative, set it equal to zero and solve there is probably an easier way, but this will certainly do it

OpenStudy (anonymous):

the derivative is \(\frac{2x+1}{x^2+x+1}\) note that the denominator is never zero, so we don't have to worry about that one

OpenStudy (anonymous):

set the numerator equal to zero and solve (in your head) you get \(x=-\frac{1}{2}\) also knows as the first coordinate of the vertex of \(x^2+x+1\) that will give you the min to find the max, check the right hand endpoints of the interval

OpenStudy (anonymous):

that is, compute \(f(-1),f(1)\) and \(f(\frac{1}{2})\) the smallest is the min, the largest is the max

OpenStudy (anonymous):

oops typo there i meant of course check \(f(-\frac{1}{2})\)

OpenStudy (anonymous):

oh okay, i got it now.

OpenStudy (anonymous):

good!